OpenStudy (anonymous):

Find the set of 4 consecutive even integers whose sum is 44.

OpenStudy (anonymous):

Hint: Call the first integer in the set 2n.

OpenStudy (anonymous):

would 3x+6=44 work?

OpenStudy (anonymous):

Can you explain your thought process? =)

OpenStudy (anonymous):

x + (x+2) + (x+4) + (x + 6) = 44

OpenStudy (anonymous):

so will 4x+12=44 work

OpenStudy (anonymous):

would it =8, 10, 12, 14

OpenStudy (anonymous):

Yeah :)